OpenStudy (anonymous):

Triangle ABC has side lengths 7, 24, and 25. Do the side lengths form a Pythagorean triple? Explain.

OpenStudy (anonymous):

since 25 is the bigger - it is the hypothen to determine this find a^2 + b^2

OpenStudy (anonymous):

7^2 + 24^2 = 25^ = if they are equal = yes they are triples if c^2 > a^2 + b^2 then this is obtuse if c^2 < a^2 + b^2 this is acute

OpenStudy (anonymous):

let sides of triangle be a , b and c be of length 7cm,24cm 1nd 25cm. a+b>c, then this triangle can be formed let us use pythagorus theorem a^2 + b^2 = c^2 7^2 + 24^2 49 + 576 625=c^2 c= 25 there fore triangle ABC is a right angled triangle